# Language steering in latent space to mitigate unintended code-switching

### v1

@@ 5 Conclusion
We introduced latent-space language steering, a lightweight inference-time method that mitigates unintended code-switching in multilingual LLMs without fine-tuning. By extracting language directions via PCA on parallel translations, our approach enables precise control over language identity with negligible overhead—one dot product and vector subtraction per token.
@@
Our empirical analysis across Qwen2.5 and Llama-3.2 models reveals that language identity concentrates in final layers with near-perfect linear separability. This geometric structure enables both accurate language classification (95-99% accuracy) and effective steering that manages to restore the original language and reduces distributional divergence by up to 42% while maintaining semantic coherence. The layer-wise evolution shows language identity crystallizes in final layers while early layers focus on semantics, suggesting these occupy largely orthogonal subspaces.
While performance is strongest for typologically distant pairs (English-Chinese), the 20% average divergence reduction indicates that simple linear transformations cannot fully reconstruct monolingual distributions. Our approach demonstrates that interpretable control mechanisms can be extracted from learned representations, offering a practical alternative to expensive fine-tuning for production multilingual systems.
@@
  • Incomplete reconstruction. Average 20% KL divergence reduction indicates linear projections cannot fully reconstruct monolingual distributions, suggesting non-linear language-semantic interactions at the manifold level.
  • Dataset coverage. Evaluation relies on Flores Plus and TED Talks (formal text). Conversational data, technical domains, and naturally occurring code-switching patterns are not tested. Artificial code-switching setup may not reflect real-world scenarios.
  • Language typology. Limited to Indo-European and Chinese languages. Low-resource languages, right-to-left scripts, and typologically distant languages (agglutinative, tonal) unexplored. Hindi showed no improvement, suggesting tokenization or data balance issues.
@@
  • Evaluation metrics. Primary reliance on KL divergence measures statistical similarity but not semantic preservation, fluency, or downstream task performance. Human evaluation and task-specific metrics needed.
@@
  • Steering calibration. Optimal coefficients require manual grid search per language pair. Principled or adaptive selection methods not developed.
@@
  • Efficiency benchmarking. No rigorous latency profiling across hardware configurations or batch sizes to validate negligible overhead claims for production deployment.

### v3

@@ 5. Conclusion
We introduced latent-space language steering, a lightweight inference-time method that mitigates unintended code-switching in multilingual LLMs without fine-tuning. By extracting language directions via PCA on parallel translations, our approach enables precise control over language identity with negligible overhead - one dot product and vector subtraction per token.
@@
Our empirical analysis across Qwen2.5 and Llama-3.2 models reveals that language identity concentrates in final layers with near-perfect linear separability. This geometric structure enables both accurate language classification (95-99% accuracy) and effective steering that reduces distributional divergence by up to 55% and achieves 63–99% Code-Switching Index reduction in generated text. The layer-wise evolution shows language identity crystallizes in final layers while early layers focus on semantics, suggesting these occupy largely orthogonal subspaces.
While distributional divergence is reduced by 40% on average and generated text shows near-complete language conversion (TLC >0.77>0.77 across all pairs), steering coefficients are highly fragile - aggressive values eliminate code-switching but can produce degenerate repetition, revealing a sharp trade-off between language control and generation fluency. Our approach demonstrates that interpretable control mechanisms can be extracted from learned representations, offering a promising direction toward lightweight alternatives to fine-tuning for multilingual language control, though addressing the fluency–control trade-off and coefficient sensitivity remains important future work. Natural extensions include multilingual steering (controlling language choice among three or more languages simultaneously) and bidirectional steering (e.g., steering from English toward Spanish), both of which the geometric structure of the multilingual PCA maps suggests should be feasible.
@@
  • Incomplete reconstruction. While the method achieves up to 55% KL divergence reduction, the remaining gap indicates linear projections cannot fully reconstruct monolingual distributions, suggesting non-linear language-semantic interactions at the manifold level.
  • Dataset coverage. Evaluation relies on Flores Plus and TED Talks (formal text). Conversational data, technical domains, and naturally occurring code-switching patterns are not tested. Our artificial code-switching setup produces mid-sentence switches at fixed positions, which may not reflect the morphosyntactic boundaries where natural code-switching typically occurs. Evaluation on naturalistic code-switching corpora such as the LinCE benchmark [1] is an important direction for future work.
  • Language typology. Limited to Indo-European and Chinese languages. Low-resource languages, right-to-left scripts, and typologically distant languages (agglutinative, tonal) unexplored. Hindi showed the weakest improvement across both models, with KL reduction (24-27%) but minimal gains in other distributional similarity metrics, suggesting tokenization or data balance issues.
@@
  • Evaluation metrics. While we supplement distributional metrics (KL, JS divergence) with generation-based code-switching metrics (CSI, M-Index, I-Index), we do not evaluate semantic preservation or downstream task performance. Human evaluation of output quality remains needed.
  • Fluency–control trade-off and coefficient fragility. Generation-based evaluation reveals that aggressive steering coefficients, while effective at eliminating code-switching (CSI <0.01<0.01 for English-Russian and English-Hindi), can produce repetitive or degenerate output patterns. Steering coefficients are highly fragile. Small changes in magnitude shift behavior from fluent code-switched text to degenerate repetition. The CSI-optimized coefficients cluster near the maximum tested values (5.0-5.0), suggesting that direct optimization of generation metrics pushes toward an aggressive regime. Combining steering with constrained decoding or adaptive coefficient scheduling could mitigate this issue.
@@
  • Steering calibration. Optimal coefficients require manual grid search per language pair and depend on the optimization target (distributional vs. generation-based). Principled or adaptive selection methods not developed.
@@
  • Efficiency benchmarking. No rigorous latency profiling across hardware configurations or batch sizes to validate overhead claims for production deployment.
  • Bilingual, English-only steering. Our evaluation steers exclusively toward English in bilingual settings. Multilingual steering (controlling language choice among three or more languages simultaneously) and bidirectional steering (e.g., steering from English toward Spanish) are natural extensions that we leave to future work. The multilingual PCA maps (Figures 2(a), 2(b)) suggest the geometric structure supports such extensions, as multiple language clusters are simultaneously separable.
Acknowledgements
The research was supported by the Russian Science Foundation grant No. 25-11-00355, https://rscf.ru/project/25-11-00355/.
Список литературы
  • [1] G. Aguilar, S. Kar, and T. Solorio (2020-05) LinCE: a centralized benchmark for linguistic code-switching evaluation. In Proceedings of the Twelfth Language Resources and Evaluation Conference, N. Calzolari, F. Béchet, P. Blache, K. Choukri, C. Cieri, T. Declerck, S. Goggi, H. Isahara, B. Maegaard, J. Mariani, H. Mazo, A. Moreno, J. Odijk, and S. Piperidis (Eds.), Marseille, France, pp. 1803–1813. External Links: Link Cited by: §2, 3rd item.
  • [2] K. Bali, J. Sharma, M. Choudhury, and Y. Vyas (2014-10) ‘‘i am borrowing ya mixing ?’’ an analysis of English-Hindi code mixing in Facebook. In Proceedings of the First Workshop on Computational Approaches to Code Switching, M. Diab, J. Hirschberg, P. Fung, and T. Solorio (Eds.), Doha, Qatar, pp. 116–126. External Links: Link, Document Cited by: §2.
  • [3] N. Belrose, D. Schneider-Joseph, S. Ravfogel, R. Cotterell, E. Raff, and S. Biderman (2023) LEACE: perfect linear concept erasure in closed form. In Advances in Neural Information Processing Systems, Vol. 36. External Links: Link Cited by: §2.
  • [4] E. A. Chi, J. Hewitt, and C. D. Manning (2020-07) Finding universal grammatical relations in multilingual BERT. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, D. Jurafsky, J. Chai, N. Schluter, and J. Tetreault (Eds.), Online, pp. 5564–5577. External Links: Link, Document Cited by: §2.
  • [5] A. Conneau, S. Wu, H. Li, L. Zettlemoyer, and V. Stoyanov (2020-07) Emerging cross-lingual structure in pretrained language models. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, D. Jurafsky, J. Chai, N. Schluter, and J. Tetreault (Eds.), Online, pp. 6022–6034. External Links: Link, Document Cited by: §2, §4.1.
  • [6] B. Gambäck and A. Das (2016-05) Comparing the level of code-switching in corpora. In Proceedings of the Tenth International Conference on Language Resources and Evaluation (LREC’16), N. Calzolari, K. Choukri, T. Declerck, S. Goggi, M. Grobelnik, B. Maegaard, J. Mariani, H. Mazo, A. Moreno, J. Odijk, and S. Piperidis (Eds.), Portorož, Slovenia, pp. 1850–1855. External Links: Link Cited by: §4.4.
  • [7] A. Grattafiori, A. Dubey, A. Jauhri, A. Pandey, A. Kadian, A. Al-Dahle, et al. (2024) The llama 3 herd of models. External Links: 2407.21783, Link Cited by: §4.
  • [8] G. GuzmГғn, J. Ricard, J. Serigos, B. E. Bullock, and A. J. Toribio (2017) Metrics for Modeling Code-Switching Across Corpora. In Interspeech 2017, pp. 67–71. External Links: Document, ISSN 2958-1796 Cited by: §4.4.
  • [9] K. Li, O. Patel, F. Viégas, H. Pfister, and M. Wattenberg (2024) Inference-time intervention: eliciting truthful answers from a language model. In Advances in Neural Information Processing Systems, Vol. 36. External Links: Link Cited by: §2.
  • [10] NLLB Team, M. R. Costa-jussà, J. Cross, O. Çelebi, M. Elbayad, K. Heafield, et al. (2024) Scaling neural machine translation to 200 languages. Nature 630 (8018), pp. 841–846. External Links: ISSN 1476-4687, Document, Link Cited by: §4.1, §4.2.
  • [11] Qwen, :, A. Yang, B. Yang, B. Zhang, B. Hui, et al. (2025) Qwen2.5 technical report. External Links: 2412.15115, Link Cited by: §4.
  • [12] S. Ravfogel, Y. Elazar, H. Gonen, M. Twiton, and Y. Goldberg (2020-07) Null it out: guarding protected attributes by iterative nullspace projection. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, D. Jurafsky, J. Chai, N. Schluter, and J. Tetreault (Eds.), Online, pp. 7237–7256. External Links: Link, Document Cited by: §2.
  • [13] M. J. Ryan, W. Held, and D. Yang (2024-08) Unintended impacts of LLM alignment on global representation. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Bangkok, Thailand, pp. 16121–16140. External Links: Link Cited by: §1, §2.
  • [14] A. M. Turner, L. Thiergart, G. Leech, D. Udell, J. J. Vazquez, U. Mini, and M. MacDiarmid (2024) Steering language models with activation engineering. External Links: 2308.10248, Link Cited by: §2.
  • [15] C. Wendler, V. Veselovsky, G. Monea, and R. West (2024-08) Do llamas work in English? on the latent language of multilingual transformers. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), L. Ku, A. Martins, and V. Srikumar (Eds.), Bangkok, Thailand, pp. 15366–15394. External Links: Link Cited by: §2, §4.1.
  • [16] Z. Yang, Y. Yang, D. Cer, and E. Darve (2021-11) A simple and effective method to eliminate the self language bias in multilingual representations. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, Online and Punta Cana, Dominican Republic, pp. 5825–5832. External Links: Link Cited by: §2.
  • [17] H. Yoo, Y. Yang, and H. Lee (2024) Code-switching red-teaming: LLM evaluation for safety and multilingual understanding. arXiv preprint arXiv:2406.15481. External Links: Link Cited by: §1, §2.
  • [18] A. Zou, L. Phan, S. Chen, J. Campbell, P. Guo, R. Ren, A. Pan, X. Yin, M. Mazeika, A. Dombrowski, S. Goel, N. Li, M. J. Byun, Z. Wang, A. Mallen, S. Basart, S. Koyejo, D. Song, M. Fredrikson, J. Z. Kolter, and D. Hendrycks (2023) Representation engineering: a top-down approach to ai transparency. External Links: 2310.01405, Link Cited by: §2.